OpenStudy (campayne):

Assume triangle JKL is in the first quadrant, with the measure of angle K = 90°. Suppose triangle JKL is a 45°-45°-90° triangle and segment JK is one of the legs. What are the coordinates of point L? (6, 2) (7, 2) (-2, 2) (-3, 2) Assume triangle JKL is in the first quadrant, with the measure of angle K = 90°. Suppose triangle JKL is a 30°-60°-90° triangle and segment JK is the side opposite the 60° angle. What are the approximate coordinates of point L? (4.5, 2) (8.7, 2) (7.1, 2) (4.9, 2)

OpenStudy (campayne):

first question here

OpenStudy (campayne):

second one

OpenStudy (evoker):

For a 45, 45,90 triangle the two legs are equal in length, for a 30, 60, 90 triangle, I would suggest finding the length and seeing which has the appropriate sin and cos.

OpenStudy (campayne):

but how???????

OpenStudy (campayne):

@sammixboo

OpenStudy (evoker):

Ok well you can find the length for the first problem, which is 5. so for the first it will be to the right by 5 or 7,2.

OpenStudy (evoker):

For the second case perhaps use tan angle = o / a. so tan(60)=5/a, therefore a = 5/tan(60) or a= 2.9 , so 2.9 to the right or 4.9, 2